# SimVG: A Simple Framework for Visual Grounding with Decoupled Multi-modal Fusion

### v1

@@ Abstract
Visual grounding is a common vision task that involves grounding descriptive sentences to the corresponding regions of an image. Most existing methods use independent image-text encoding and apply complex hand-crafted modules or encoder-decoder architectures for modal interaction and query reasoning. However, their performance significantly drops when dealing with complex textual expressions. This is because the former paradigm only utilizes limited downstream data to fit the multi-modal feature fusion. Therefore, it is only effective when the textual expressions are relatively simple. In contrast, given the wide diversity of textual expressions and the uniqueness of downstream training data, the existing fusion module, which extracts multimodal content from a visual-linguistic context, has not been fully investigated. In this paper, we present a simple yet robust transformer-based framework, SimVG, for visual grounding. Specifically, we decouple visual-linguistic feature fusion from downstream tasks by leveraging existing multimodal pre-trained models and incorporating additional object tokens to facilitate deep integration of downstream and pre-training tasks. Furthermore, we design a dynamic weight-balance distillation method in the multi-branch synchronous learning process to enhance the representation capability of the simpler branch. This branch only consists of a lightweight MLP, which simplifies the structure and improves reasoning speed. Experiments on six widely used VG datasets, i.e., RefCOCO/+/g, ReferIt, Flickr30K, and GRefCOCO, demonstrate the superiority of SimVG. Finally, the proposed method not only achieves improvements in efficiency and convergence speed but also attains new state-of-the-art performance on these benchmarks. Codes and models will be available at https://github.com/Dmmm1997/SimVG.
@@ 1 Introduction
Visual grounding (VG) aims to predict the corresponding regions of an image through linguistic expressions. The task necessitates a comprehensive understanding of each modality, as well as the modeling of consistency between image context and text. Some benchmarks focus on addressing phrase localization [24, 48], which entails locating all objects mentioned in a sentence within an image. Another aspect emphasizes resolving referring expression comprehension (REC) [73, 46, 45], characterized by only one target corresponding to a sentence. Recently, a new type of general referring expression comprehension (GREC) [35, 16] task has emerged. GREC is similar to REC, but in which a sentence can have multiple targets or no target at all.
Existing visual grounding models can be roughly divided into three categories: two-stage, one-stage, and transformer-based. Among them, as shown in Fig. 1(a), two-stage methods [18, 75, 72, 42, 36] require a pre-trained detector to generate proposals and perform localization through region-text retrieval. These methods rely on a complex module with manually designed mechanisms to achieve query reasoning and multi-modal fusion. One-stage methods [69, 34, 44, 67], on the other hand, employ an end-to-end architecture, as shown in Fig. 1(b). Most of them primarily perform dense prediction on multimodal fusion features defined in the form of anchors. Some recent algorithms [7, 23, 79, 54], depicted in Fig. 1(c), adopt an encoder-decoder architecture to perform multimodal fusion in the encoder and then decode the response target position using an object query similar to DETR[2]. The existing methods share a commonality: they adopt architectures that independently encode each modality before merging them, with multimodal fusion intricately linked to each visual grounding task. The feature extraction part of these methods generally employs specific classification [6, 53] or autoregressive [8, 29] tasks in each modality for pre-training. However, the alignment and mutual understanding between modalities only utilize a limited amount of downstream data, which undoubtedly underestimates the difficulty of achieving mutual understanding between modalities. Another observed trend, as noted in [23, 79], is the notable enhancement in the performance of visual grounding with a significant augmentation of pretraining data on large corpora. This implicitly suggests that leveraging a small amount of downstream data does not fully capitalize on the potential for mutual understanding between images and text. Nevertheless, this type of pretraining undoubtedly increases the burden of training resources.
@@
To be specific, we delve into existing multimodal understanding methods in the form of multimodal pre-training. Existing approaches can be broadly categorized into three categories. Dual-stream structures [49, 22, 32] encode image-text modalities independently and supervise them with contrastive learning. One-stream models [5, 25, 29, 21] concatenate multimodal features for feature extraction. Other works [30, 1, 55, 60] use a dual-stream design with a fusion encoder to balance complexity and computational cost, fusing multimodal features in intermediate layers. Some methods also have applied multimodal pre-training to VG tasks, such as Dynamic MDETR [54], which uses an image-text encoder pretrained on CLIP [49] to enhance the performance. Recently, works like CPT [70], ReCLIP [56], and FGVP [65] have improved the performance of zero-shot visual grounding using dual-stream pre-trained models that employ a two-stage-like architecture and prompt engineering techniques. However, these approaches focus on multimodal alignment rather than mutual understanding. Thus, the dual-stream with fusion encoder architecture [30, 1, 55, 60] has been thrown into our sight. Specifically, based on BEiT-3 [60], we propose a simple framework called SimVG that decouples multimodal fusion from downstream tasks and simultaneously encodes image, text, and object tokens, as illustrated in Fig. 1 (d).
@@
The decoder structure used for query reasoning, similar to DETR [2], is effective but inevitably increases the model’s complexity and computational overhead. We aim to develop a more efficient and simpler architecture for visual grounding. To achieve this, in addition to the decoder branch, we introduce a lightweight token branch. This branch leverages object tokens that are deeply integrated with image-text features to enhance grounding performance. To ensure that the token branch maintains high performance while enabling efficient inference, we adhere to the principles of the existing knowledge distillation methods [20, 4, 3] and introduce an innovative dynamic weight-balance distillation (DWBD). This approach enhances the token branch’s capability by dynamically assigning weights to the decoder’s predictions and ground truth at various learning stages, facilitating more effective learning in the token branch. Furthermore, we introduce a text-guided query generation (TQG) module to integrate textual information into queries, enabling the adaptive incorporation of textual prior knowledge into object queries. Notably, the design of the TQG module allows for the expansion of object tokens, increasing the number of objects it can handle and thereby adapting effectively to the GREC [16] task. The experiments demonstrate that, by decoupling multimodal fusion from downstream tasks, SimVG achieves rapid convergence and superior performance even with a limited amount of data. Coupled with our proposed DWBD and TQG modules, SimVG sets new state-of-the-art performance across various benchmarks.
@@ 2 Related Work
Existing vision-language pretraining (VLP) models can be broadly categorized into three main types: one-stream, dual-stream, and dual-stream with fusion encoder architectures. One-stream models [5, 25, 29, 21] process both image and text inputs in a single stream, concatenate image and text embeddings, and interact cross-modals information throughout the whole feature extraction process. In contrast, dual-stream models [49, 22, 32] employ separate encoders for each modality. These models do not concatenate modalities at the input level. Instead, the interaction between pooled image and text vectors occurs at a shallow layer. Dual-stream models with fusion encoder [30, 1, 55, 60] combine aspects of both one-stream and dual-stream models. They allow for intermediate interaction between modalities, potentially offering a balance between complexity and performance. In this paper, we improve the performance of visual grounding by decoupling multi-modal fusion from downstream tasks into upstream VLP models [60].
@@
Early approaches in REC typically followed a two-stage pipeline [18, 75, 80, 72, 42, 66, 17, 36]. This pipeline involves first extracting region proposals [51], which are then ranked based on their similarity scores with the language query. In contrast, a more recent line of research [77, 69, 34, 44, 67, 19] advocates for a simpler and faster one-stage pipeline based on dense anchors. Several recent approaches [7, 78, 23, 9, 31, 79] have employed a transformer-based structure [58] for multi-modal fusion. Furthermore, with the vigorous development of multimodal large language models (MLLM) [71], some of the latest methods have further enhanced the generalization performance of REC through zero-shot [65, 56] or fine-tuning [63, 47] methods, leveraging the powerful capabilities of large models [37] and general models [27]. In contrast to existing methods, our proposed SimVG method directly feeds object, image, and text tokens into the multi-modality encoder for multimodal feature interaction. We eschew the complex encoder-decoder structure and perform visual grounding directly using a simple MLP.
@@
The majority of research in knowledge distillation has primarily focused on classification tasks [14, 62, 74]. Several studies [43, 61, 76, 59] have extended knowledge distillation techniques to dense prediction tasks, such as semantic segmentation and object detection. These works commonly exploit pixel-wise correlations or channel-wise interactions between dense features of teacher and student models. Recently, there has been a growing interest in developing tailored knowledge distillation losses for query-based detectors like DETR [2], as demonstrated by works such as [20, 4, 3]. Unlike previous methods that guide a lightweight student model using a pre-trained teacher model, this paper introduces knowledge distillation during synchronous learning to enhance the performance of the lightweight branch.
Figure 3: Overview of the proposed SimVG network with TQG module structure and DWBD process illustration. The token branch refers to the upper light yellow region, while the decoder branch refers to the lower light blue region. During model inference, we can independently apply the more lightweight token branch to improve inference speed and simplify the model architecture.
@@ 3.3 Text-guided Query Generation
The initial object queries, $O_{init}$, are defined using learnable embeddings, without any prior information to guide them. From a macro perspective, visual grounding involves using text as a query to locate optimal regions in an image. Embedding text into queries to adaptively provide priors offers a viable solution. Therefore, we propose a text-guided query generation (TQG) module to generate object queries with text priors. As illustrated in Fig. 3, the process of generating queries through TQG can be expressed as follows:
@@ 4.2 Implementation Details
We train SimVG 30 epochs for REC and phrase localization, and 200 epochs for GREC, using a batch size of 32. The Adam [26] optimizer with an initial learning rate of 5e-4 is employed, with a 10 times decay after 25 epochs. Following standard practices, images are resized to 640$\times$640, and the length of language expressions is trimmed to 20 for all the datasets. For pre-training, SimVG is trained for 30 epochs and then fine-tuned for another 10 epochs. The pre-training experiments are run on 8 NVIDIA RTX 3090 GPUs. All the other experiments are conducted on 2 NVIDIA RTX 4090 GPUs. More implementation details are reported in Appendix C.
@@ 4.3 Comparison with The State-of-the-Art
In this part, we compare our SimVG with the SOTA methods on six mainstream datasets, i.e., RefCOCO/+/g, GRefCOCO, ReferItGame, and Flickr30K. We combine the results of RefCOCO/+/g, ReferItGame and Flickr30K datasets in Table 1, and the results of GREC are reported in Table 2. Table 3 reports the results of the methods pre-trained on the large corpus of data.
@@
As shown in Table 1, on the ReferItGame and Flickr30K Entities datasets which mostly contain short noun phrases, the performance boosts to 74.83 and 82.04 with a large margin over the previous one-stage method [67]. Compared to existing transformer-based methods [7, 31, 79], SimVG still significantly outperforms most SOTA methods by approximately 3.4 points on the ReferItGame dataset, and it also slightly outperforms Dynamic MDETR [54] on the Flickr30k dataset. Furthermore, scaling the model from base to large has led to significant improvements across all the datasets.
@@
Table 3 demonstrates that when pre-training on a large corpus of image-text pairs, SimVG exhibits greater data efficiency as compared with most of the existing SOTA methods. Despite utilizing only 28K images, which is nearly six times fewer than MDETR [23], and three times fewer than RefTR [31], SimVG still achieves SOTA performance, surpassing most existing methods by a significant margin. Compared to MDETR, SimVG demonstrates an average improvement of 5 points, and compared to the recent SOTA model GroundingDINO [41], it achieves an average improvement of 2 points. Moreover, increasing the volume of pre-training data further enhances performance. Additionally, SimVG applies a lighter transformer structure in the head. Specifically, SimVG-TB only uses 1.58 million parameters, which is smaller than some lightweight models [79, 31]. Lastly, we observe that scaling the multimodal encoder from ViT-B to ViT-L results in the performance of the lightweight token branch surpassing that of its teacher model. We hypothesize that as the model size increases, the reliability of the decoder branch’s performance improves, helping to mitigate the impact of mislabeled ground truth data. This, in turn, enhances the generalization ability of the token branch, further demonstrating the effectiveness of the DWBD method.
@@ 4.4 Ablation Studies
Table 4: Some ablation experiments on different multimodal fusion architectures. VE Interp. refers to the downsampling convolution kernel in Visual Embed that performs bilinear interpolation from pre-trained weights.
@@ 4.4.1 Multi-Modality Encoder Architecture
Table 5: Ablation studies related to TQG module in the token and decoder branches.
@@ 4.4.2 Text-guided Query Generation
Table 6: Ablation studies related to DWBD module, including one/two stage distillation results.
@@ D.1 Number of Layers and Query
The number of transformer layers in the decoder branch and MLP layers in the token branch can impact performance and efficiency. Excessive layers can reduce computational efficiency and increase parameter count, while too few layers may lead to subpar performance. To determine the optimal number of layers, we conducted experiments under varying settings, with results presented in Fig. 7 and Fig. 8. Fig. 7 indicates that increasing the number of MLP layers does not yield significant gains. Therefore, we selected single MLP layer. In contrast, Fig. 8 shows that increasing the number of transformer layers results in performance improvement. Considering efficiency, we opted for 3 transformer layers.
@@
D.2 Compare with Multi-modal Large Language Models

### v2

@@ Abstract
Visual grounding is a common vision task that involves grounding descriptive sentences to the corresponding regions of an image. Most existing methods use independent image-text encoding and apply complex hand-crafted modules or encoder-decoder architectures for modal interaction and query reasoning. However, their performance significantly drops when dealing with complex textual expressions. This is because the former paradigm only utilizes limited downstream data to fit the multi-modal feature fusion. Therefore, it is only effective when the textual expressions are relatively simple. In contrast, given the wide diversity of textual expressions and the uniqueness of downstream training data, the existing fusion module, which extracts multimodal content from a visual-linguistic context, has not been fully investigated. In this paper, we present a simple yet robust transformer-based framework, SimVG, for visual grounding. Specifically, we decouple visual-linguistic feature fusion from downstream tasks by leveraging existing multimodal pre-trained models and incorporating additional object tokens to facilitate deep integration of downstream and pre-training tasks. Furthermore, we design a dynamic weight-balance distillation method in the multi-branch synchronous learning process to enhance the representation capability of the simpler branch. This branch only consists of a lightweight MLP, which simplifies the structure and improves reasoning speed. Experiments on six widely used VG datasets, i.e., RefCOCO/+/g, ReferIt, Flickr30K, and GRefCOCO, demonstrate the superiority of SimVG. Finally, the proposed method not only achieves improvements in efficiency and convergence speed but also attains new state-of-the-art performance on these benchmarks. Codes and models are available at https://github.com/Dmmm1997/SimVG.
@@ 1 Introduction
Visual grounding (VG) aims to predict the corresponding regions of an image through linguistic expressions. The task necessitates a comprehensive understanding of each modality, as well as the modeling of consistency between image context and text. Some benchmarks focus on addressing phrase localization [26; 50], which entails locating all objects mentioned in a sentence within an image. Another aspect emphasizes resolving referring expression comprehension (REC) [80; 48; 47], characterized by only one target corresponding to a sentence. Recently, a new type of general referring expression comprehension (GREC) [37; 17] task has emerged. GREC is similar to REC, but in which a sentence can have multiple targets or no target at all.
Existing visual grounding models can be roughly divided into three categories: two-stage, one-stage, and transformer-based. Among them, as shown in Fig. 1(a), two-stage methods [19; 83; 79; 44; 38] require a pre-trained detector to generate proposals and perform localization through region-text retrieval. These methods rely on a complex module with manually designed mechanisms to achieve query reasoning and multi-modal fusion. One-stage methods [75; 36; 46; 73], on the other hand, employ an end-to-end architecture, as shown in Fig. 1(b). Most of them primarily perform dense prediction on multimodal fusion features defined in the form of anchors. Some recent algorithms [8; 25; 87; 57], depicted in Fig. 1(c), adopt an encoder-decoder architecture to perform multimodal fusion in the encoder and then decode the response target position using an object query similar to DETR[2]. The existing methods share a commonality: they adopt architectures that independently encode each modality before merging them, with multimodal fusion intricately linked to each visual grounding task. The feature extraction part of these methods generally employs specific classification [7; 55] or autoregressive [9; 30] tasks in each modality for pre-training. However, the alignment and mutual understanding between modalities only utilize a limited amount of downstream data, which undoubtedly underestimates the difficulty of achieving mutual understanding between modalities. Another observed trend, as noted in [25; 87], is the notable enhancement in the performance of visual grounding with a significant augmentation of pretraining data on large corpora. This implicitly suggests that leveraging a small amount of downstream data does not fully capitalize on the potential for mutual understanding between images and text. Nevertheless, this type of pretraining undoubtedly increases the burden of training resources.
@@
To be specific, we delve into existing multimodal understanding methods in the form of multimodal pre-training. Existing approaches can be broadly categorized into three categories. Dual-stream structures [51; 23; 33] encode image-text modalities independently and supervise them with contrastive learning. One-stream models [6; 27; 30; 22] concatenate multimodal features for feature extraction. Other works [31; 1; 58; 64] use a dual-stream design with a fusion encoder to balance complexity and computational cost, fusing multimodal features in intermediate layers. Some methods also have applied multimodal pre-training to VG tasks, such as Dynamic MDETR [57], which uses an image-text encoder pretrained on CLIP [51] to enhance the performance. Recently, works like CPT [76], ReCLIP [59], and FGVP [71] have improved the performance of zero-shot visual grounding using dual-stream pre-trained models that employ a two-stage-like architecture and prompt engineering techniques. However, these approaches focus on multimodal alignment rather than mutual understanding. Thus, the dual-stream with fusion encoder architecture [31; 1; 58; 64] has been thrown into our sight. Specifically, based on BEiT-3 [64], we propose a simple framework called SimVG that decouples multimodal fusion from downstream tasks and simultaneously encodes image, text, and object tokens, as illustrated in Fig. 1 (d). The decoder structure used for query reasoning, similar to DETR [2], is effective but inevitably increases the model’s complexity and computational overhead. We aim to develop a more efficient and simpler architecture for visual grounding. To achieve this, in addition to the decoder branch, we introduce a lightweight token branch. This branch leverages object tokens that are deeply integrated with image-text features to enhance grounding performance. To ensure that the token branch maintains high performance while enabling efficient inference, we adhere to the principles of the existing knowledge distillation methods [21; 5; 3] and introduce an innovative dynamic weight-balance distillation (DWBD). This approach enhances the token branch’s capability by dynamically assigning weights to the decoder’s predictions and ground truth at various learning stages, facilitating more effective learning in the token branch. Furthermore, we introduce a text-guided query generation (TQG) module to integrate textual information into queries, enabling the adaptive incorporation of textual prior knowledge into object queries. Notably, the design of the TQG module allows for the expansion of object tokens, increasing the number of objects it can handle and thereby adapting effectively to the GREC [17] task. The experiments demonstrate that, by decoupling multimodal fusion from downstream tasks, SimVG achieves rapid convergence and superior performance even with a limited amount of data. Coupled with our proposed DWBD and TQG modules, SimVG sets new state-of-the-art performance across various benchmarks.
@@ 2 Related Work
Existing vision-language pretraining (VLP) models can be broadly categorized into three main types: one-stream, dual-stream, and dual-stream with fusion encoder architectures. One-stream models [6; 27; 30; 22] process both image and text inputs in a single stream, concatenate image and text embeddings, and interact cross-modals information throughout the whole feature extraction process. In contrast, dual-stream models [51; 23; 33] employ separate encoders for each modality. These models do not concatenate modalities at the input level. Instead, the interaction between pooled image and text vectors occurs at a shallow layer. Dual-stream models with fusion encoder [31; 1; 58; 64] combine aspects of both one-stream and dual-stream models. They allow for intermediate interaction between modalities, potentially offering a balance between complexity and performance. In this paper, we improve the performance of visual grounding by decoupling multi-modal fusion from downstream tasks into upstream VLP models [64].
@@
Early approaches in REC typically followed a two-stage pipeline [19; 83; 88; 79; 44; 72; 18; 38]. This pipeline involves first extracting region proposals [53], which are then ranked based on their similarity scores with the language query. In contrast, a more recent line of research [85; 75; 36; 46; 73; 20] advocates for a simpler and faster one-stage pipeline based on dense anchors. Several recent approaches [8; 86; 25; 10; 32; 87] have employed a transformer-based structure [61] for multi-modal fusion. Furthermore, with the vigorous development of multimodal large language models (MLLM) [77], some of the latest methods have further enhanced the generalization performance of REC through zero-shot [71; 59] or fine-tuning [69; 49] methods, leveraging the powerful capabilities of large models [39] and general models [28]. In contrast to existing methods, our proposed SimVG method directly feeds object, image, and text tokens into the multi-modality encoder for multimodal feature interaction. We eschew the complex encoder-decoder structure and perform visual grounding directly using a simple MLP.
@@
The majority of research in knowledge distillation has primarily focused on classification tasks [15; 66; 81]. Several studies [45; 65; 84; 63] have extended knowledge distillation techniques to dense prediction tasks, such as semantic segmentation and object detection. These works commonly exploit pixel-wise correlations or channel-wise interactions between dense features of teacher and student models. Recently, there has been a growing interest in developing tailored knowledge distillation losses for query-based detectors like DETR [2], as demonstrated by works such as [21; 5; 3]. Unlike previous methods that guide a lightweight student model using a pre-trained teacher model, this paper introduces knowledge distillation during synchronous learning to enhance the performance of the lightweight branch.
Figure 3: Overview of the proposed SimVG. The token branch refers to the upper light yellow region, while the decoder branch refers to the lower light blue region. During model inference, we can independently apply the more lightweight token branch to improve inference speed and simplify the model architecture.
@@ 3.3 Text-guided Query Generation
The initial object queries, $Q_{init}$, are defined using learnable embeddings, without any prior information to guide them. From a macro perspective, visual grounding involves using text as a query to locate optimal regions in an image. Embedding text into queries to adaptively provide priors offers a viable solution. Therefore, we propose a text-guided query generation (TQG) module to generate object queries with text priors. As illustrated in Fig. 3, the process of generating queries through TQG can be expressed as follows:
@@ 4.2 Implementation Details
We train SimVG 30 epochs for REC and phrase localization, and 200 epochs for GREC, using a batch size of 32. Following standard practices, images are resized to 640$\times$640, and the length of language expressions is trimmed to 20 for all the datasets. For pre-training, SimVG is trained for 30 epochs and then fine-tuned for another 10 epochs. The pre-training experiments are run on 8 NVIDIA RTX 3090 GPUs. All the other experiments are conducted on 2 NVIDIA RTX 4090 GPUs. More implementation details are reported in Appendix C.
@@ 4.3 Comparison with The State-of-the-art
In this part, we compare our SimVG with the SOTA methods on six mainstream datasets. We combine the results of RefCOCO/+/g, ReferItGame and Flickr30K datasets in Table 1, and the results of GREC are reported in Table 2. Table 3 reports the results pre-trained on the large corpus of data.
@@
As shown in Table 1, on the ReferItGame and Flickr30K Entities datasets which mostly contain short noun phrases, the performance boosts to 74.83 and 82.04 with a large margin over the previous one-stage method [73]. Compared to existing transformer-based methods [8; 32; 87], SimVG still significantly outperforms most SOTA methods by approximately 3.4 points on the ReferItGame dataset, and it also slightly outperforms Dynamic MDETR [57] on the Flickr30k dataset. Furthermore, scaling the model from base to large has led to significant improvements across all the datasets.
@@
Table 3 demonstrates that when pre-training on a large corpus of image-text pairs, SimVG exhibits greater data efficiency as compared with most of the existing SOTA methods. Despite utilizing only 28K images, which is nearly six times fewer than MDETR [25], and three times fewer than RefTR [32], SimVG still achieves SOTA performance, surpassing most existing methods by a significant margin. Compared to MDETR, SimVG demonstrates an average improvement of 5 points, and compared to the recent SOTA model GroundingDINO [43], it achieves an average improvement of 2 points. Moreover, increasing the volume of pre-training data further enhances performance. Additionally, SimVG applies a lighter transformer structure in the head. Specifically, SimVG-TB only uses 1.58 million parameters, which is smaller than some lightweight models [87; 32]. Lastly, we observe that scaling the multimodal encoder from ViT-B to ViT-L results in the performance of the lightweight token branch surpassing that of its teacher model. We hypothesize that as the model size increases, the reliability of the decoder branch’s performance improves, helping to mitigate the impact of mislabeled ground truth data. This, in turn, enhances the generalization ability of the token branch, further demonstrating the effectiveness of the DWBD method.
@@ 4.4 Ablation Studies
Figure 4: Some ablation experiments on different multimodal fusion architectures. VE Interp. refers to the downsampling convolution kernel in Visual Embed that performs bilinear interpolation from pre-trained weights.
@@ 4.4.1 Multi-Modality Encoder Architecture
Figure 6: Ablation studies related to TQG module in the token and decoder branches.
@@ 4.4.2 Text-guided Query Generation
Figure 8: Ablation studies related to DWBD module, including one/two stage distillation results.
@@ 4.4.3 Dynamic Weight-Balance Distillation
5 Visualization
Figure 10: Visualization of the feature maps in MME and the decoder. The heatmap in MME is generated using GradCAM [56], while the attention response of the decoder is obtained from the attention map.
We conduct an attention analysis of SimVG from two perspectives, as shown in Fig. 10. First, we visualize the multimodal representations of BEiT-3 using GradCAM [56] to generate heatmaps, revealing that BEiT-3 primarily focuses on global foreground information. Additionally, we visualize the attention maps of the decoder, which highlight the model’s focus on regions referred to by the text. More qualitative results can be found in Appendix G.
@@ 6 Conclusion
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grants 62276061.
@@ D.1 Number of Layers and Query
The number of transformer layers in the decoder branch and MLP layers in the token branch can impact performance and efficiency. Excessive layers can reduce computational efficiency and increase parameter count, while too few layers may lead to subpar performance. To determine the optimal number of layers, we conducted experiments under varying settings, with results presented in Fig. D.1 and Fig. 12. Fig. D.1 indicates that increasing the number of MLP layers does not yield significant gains. Therefore, we selected single MLP layer. In contrast, Fig. 12 shows that increasing the number of transformer layers results in performance improvement. Considering efficiency, we opted for 3 transformer layers.
@@
D.2 Efficiency of Training
We analyze the training efficiency of SimVG from two aspects. As shown in Table 14, SimVG significantly outperforms mainstream methods in terms of the required number of epochs and training time. On the RefCOCO+ dataset, SimVG only requires 30 epochs and 5.5 hours, faster than SeqTR’s 60 epochs and 9 hours, demonstrating the accelerated convergence brought by multimodal pre-training. As shown in Table 14, SimVG adopts a more lightweight head design and, despite using less pre-training data, achieves notable performance improvements.
Method Visual Epoch Training
Encoder Time
TransVG RN50 180 ~90h
VGTR RN50 120 ~50h
Dynamic MDETR ViT-B/16 90 -
SeqTR DN53 60 ~9h
SimVG ViT-B/32 30 ~5.5h
Figure 13: Training time on a single RTX 4090 GPU and the number of epochs required for convergence on the RefCOCO+ dataset.
Method Head Pretrain RefCOCO
Params (M) Images val testA testB
RefTR 17.86 100K 85.65 88.73 81.16
MDETR 17.36 200K 86.75 89.58 81.41
SeqTR 7.90 174K 87.00 90.15 83.59
SimVG-DB 6.32 28K 90.98 92.68 87.94
SimVG-TB 1.58 28K 90.18 92.41 87.21
Figure 14: A comparison of the number of parameters in the head and the number of pre-training images for some publicly available methods.
D.3 Compare with Multi-modal Large Language Models